Random samples of size 𝑛 = 2 are drawn from a finite population


consisting of the numbers 5,6,7,8, and 9.


a. Find the mean of the population πœ‡.


b. Find the standard deviation of the population 𝜎.


c. Find the mean of the sampling distribution of the sample means πœ‡π‘‹Μ….


d. Find the standard deviation of the sampling distribution of the sample


means πœŽπ‘‹Μ….


e. Verify the Central Limit theorem by:


I. Comparing πœ‡ and πœ‡π‘‹Μ….


II. Comparing 𝜎 and πœŽπ‘‹Μ….

1
Expert's answer
2022-03-28T04:58:50-0400

"a:\mu =\frac{5+6+7+8+9}{5}=7\\b:\sigma =\sqrt{\frac{\left( 5-7 \right) ^2+\left( 6-7 \right) ^2+\left( 7-7 \right) ^2+\left( 8-7 \right) ^2+\left( 9-7 \right) ^2}{5}}=\sqrt{2}\\c:All\,\,samples\\\left( 5,6 \right) ,\bar{x}=5.5\\\left( 5,7 \right) ,\bar{x}=6\\\left( 5,8 \right) ,\bar{x}=6.5\\\left( 5,9 \right) ,\bar{x}=7\\\left( 6,7 \right) ,\bar{x}=6.5\\\left( 6,8 \right) ,\bar{x}=7\\\left( 6,9 \right) ,\bar{x}=7.5\\\left( 7,8 \right) ,\bar{x}=7.5\\\left( 7,9 \right) ,\bar{x}=8\\\left( 8,9 \right) ,\bar{x}=8.5\\\mu _{\bar{x}}=\frac{5.5+6+2\cdot 6.5+2\cdot 7+2\cdot 7.5+8+8.5}{10}=7\\d: \sigma _{\bar{x}}=\sqrt{\frac{\left( 5.5-7 \right) ^2+\left( 6-7 \right) ^2+2\cdot \left( 6.5-7 \right) ^2+2\left( 7-7 \right) ^2+2\left( 7.5-7 \right) ^2+\left( 8-7 \right) ^2+\left( 8.5-7 \right) ^2}{10}}=\sqrt{0.75}\\e:\\I:\mu =\mu _{\bar{x}}\\II:\frac{\sigma}{\sqrt{2}}\sqrt{\frac{5-2}{5-1}}=\frac{\sqrt{2}}{\sqrt{2}}\sqrt{\frac{3}{4}}=\sqrt{0.75}=\sigma _{\bar{x}}\\"